A father is 30 years older than his son, in 17 years' time he will be double the son's age. What is the son's age?
Answer by mananth(16946)

son =x
father = x+30
17 years later
son = x+17
father = (x+30)+17
(x+30) +17 = 2(x+17)
x+47=2x+34
2x-x=47-34
x=13 son's age